OpenStudy (pphalke):

The balance on Ted's credit card on August 9, the billing date, was $1,298.51. for the period ending September 8, he made the following transactions. assume an APR of 3.9% I need help

OpenStudy (pphalke):

@Photon336 @peachpi @jim_thompson5910 @Anaise

jimthompson5910 (jim_thompson5910):

you forgot to post the balances

OpenStudy (pphalke):

August 14, Payment $350.00 August 22, Charge: Gasoline, $39.92 August 29, Charge: Groceries, $87.17

jimthompson5910 (jim_thompson5910):

ok thanks, one moment

jimthompson5910 (jim_thompson5910):

From Aug 9 to Aug 13, the balance is $1,298.51 This is a period of 5 days (13-9+1 = 5) From Aug 14 to Aug 21, the balance is $948.51 (1298.51-350 = 948.51) This is a period of 8 days (21-14+1 = 8) From Aug 22 to Aug 28, the balance is $988.43 (948.51+39.92 = 988.43) This is a period of 7 days (28-22+1 = 7) From Aug 29 to Sept 8, the balance is $1,075.6 (988.43+87.17 = 1075.6) This is a period of 11 days (3 days from Aug 29 to Aug 31; 8 days from Sept 1 to Sept 8) Do you agree with these statements?

jimthompson5910 (jim_thompson5910):

@pphalke let me know if you're stuck anywhere

OpenStudy (pphalke):

yes I agree to these statements

OpenStudy (pphalke):

@jim_thompson5910

OpenStudy (pphalke):

@jim_thompson5910 please help what doi do next